OpenStudy (anonymous):

Find three consecutive even integers such that their sum is 12 less than twice the smallest.

OpenStudy (texaschic101):

3 consecutive even integers.....x and x + 2, and x + 4 they equal 12 less then twice the smallest......2x - 12 x + (x + 2) + (x + 4) = 2x - 12 --- combine like terms 3x + 6 = 2x - 12 --- subtract 2x from both sides 3x - 2x + 6 = -12 --- subtract 6 from both sides 3x - 2x = -12 - 6 --- combine like terms x = -18 x + 2 = - 18 + 2 = -16 x + 4 = -18 + 4 = - 14 check...their sum is 12 less then twice the smallest -18 + (-16) + (-14) = -18 - 16 - 14 = - 48 (the sum of all three) - 48 = 2(-18) - 12 - 48 = -36 - 12 - 48 = - 48 (correct) your numbers are : -18, -16, -14